Stock A has a beta of 1.20 and is farily priced. The riskless rate of return​ (US Treasury​ Bills) is 3.2%. The market risk premium is 5.4​%. Construct a​ $100,000 portfolio of stock A and Treasury Bills that will have an expected return of 5.144%. (Rounded to the nearest​ dollar.)
Answer #1

first let us know the expected return on stock A = risk free rate+ beta *(risk premium)

=>3.2% + 1.20*(5.4%)

=>9.68%.

let the percentage of amount invested in stock A be x..

the percentage involved in treasury bill be (1-x).

weighted average expected return:

=> (x*0.0968) + (1-x)*0.032 =>0.05144.

=>0.0968x +0.032 -0.032x = 0.05144

=>0.0648x=0.01944

=>x = 0.01944/0.0648

=>x=0.30.

percentage of amount invested in stock A = x =0.30

amount to be invested = 100,000*0.30=>$30,000.

amount to be invested in tresury bill =>100,000-30,000=>$70,000.

the portfolio will be like below:

Treasury bill (70%) 70,000
stock A (30%) 30,000
total 100,000